# Unified Visual Understanding and Generation

### v1

@@ Abstract
Although existing unified models deliver strong performance on vision-language understanding and text-to-image generation, their models are limited in exploring image perception and manipulation tasks, which are urgently desired by users for wide applications. Recently, OpenAI released their powerful GPT-4o-Image model for comprehensive image perception and manipulation, achieving expressive capability and attracting community interests. By observing the performance of GPT-4o-Image in our carefully constructed experiments, we infer that GPT-4o-Image leverages features extracted by semantic encoders instead of VAE, while VAEs are considered essential components in many image manipulation models. Motivated by such inspiring observations, we present a unified generative framework named UniWorld based on semantic features provided by powerful visual-language models and contrastive semantic encoders. As a result, we build a strong unified model using only 1% amount of BAGEL’s data, which consistently outperforms BAGEL on image editing benchmarks. UniWorld also maintains competitive image understanding and generation capabilities, achieving strong performance across multiple image perception tasks. We fully open-source our models, including model weights, training & evaluation scripts, and datasets.
@@ 2 Observation
GPT-4o-Image [1] achieves impressive performance in various image tasks in a generative way, and we divide supported image tasks into two categories: Image Perception (detection, segmentation, depth prediction, etc.) and Image Manipulation (editing [22], reference style transfer [33], subject consistency generation [46, 45], etc.). It is generally believed within the community that GPT-4o-Image demonstrates the necessity of unifying understanding and generation by integrating an auto-regressive understanding module with a diffusion-based generation module [41] for addressing various image tasks with complex requirements. However, we consider that additional visual features beyond auto-regressive tokens from understanding models should be injected to maintain image information, since we fail to train an effective generation model from solely VLM outputs. Recent notable unified image manipulation approaches, exemplified by Step1X-Edit [22] and FLUX-Kontext [13], simultaneously introduce variational autoencoders (VAEs) to extract image featuresa as reference image control. Although their models perform well on individual editing tasks, our experiments show that they fail to converge when extended to multiple image perception and manipulation tasks, which indicates that the manner of additional visual feature injection remains underexplored. To explore the visual feature injection method that GPT-4o-Image utilizes, we construct two groups of experiments and obtain several key observations (see Figure 2), from which we infer that GPT-4o-Image likely employs visual features extracted by semantic encoders rather than VAEs.

### v2

@@ Abstract
Although existing unified models achieve strong performance in vision-language understanding and text-to-image generation, they remain limited in addressing image perception and manipulation—capabilities increasingly demanded in practical applications. Recently, OpenAI introduced the powerful GPT-4o-Image model, which showcases advanced capabilities in comprehensive image perception and manipulation, sparking widespread interest. Through carefully designed experiments, we observe that GPT-4o-Image likely relies on semantic encoders rather than VAEs for feature extraction, despite VAEs being commonly regarded as crucial for image manipulation tasks. Inspired by this insight, we propose UniWorld, a unified generative framework built upon semantic features extracted from powerful multimodal large language models and contrastive semantic encoders. Using only 2.7M training data, UniWorld achieves impressive performance across diverse tasks, including image understanding, generation, manipulation, and perception. We fully open-source the UniWorld framework, including model weights, training and evaluation scripts, and datasets to promote reproducibility and further research.
@@ 2 Observation
GPT-4o-Image [1] achieves impressive performance in various image tasks in a generative way, and we divide supported image tasks into two categories: Image Perception (detection, segmentation, depth prediction, etc.) and Image Manipulation (editing [26, 58, 6], reference style transfer [38], subject consistency generation [52, 51], etc.). It is generally believed within the community that GPT-4o-Image demonstrates the necessity of unifying understanding and generation by integrating an auto-regressive understanding module [60, 19, 21] with a diffusion-based generation module [47] for addressing various image tasks with complex requirements. However, we consider that additional visual features beyond auto-regressive tokens from understanding models should be injected to maintain image information, since we fail to train an effective generation model from solely VLM outputs. Recent notable unified image manipulation approaches, exemplified by Step1X-Edit [26] and FLUX-Kontext [15], simultaneously introduce variational autoencoders (VAEs) to extract image featuresa as reference image control. Although their models perform well on individual editing tasks, our experiments show that they fail to converge when extended to multiple image perception and manipulation tasks, which indicates that the manner of additional visual feature injection remains underexplored. To explore the visual feature injection method that GPT-4o-Image utilizes, we construct two groups of experiments and obtain several key observations (see Figure 2), from which we infer that GPT-4o-Image likely employs visual features extracted by semantic encoders rather than VAEs.
@@ 4.3 Image Manipulation
Table 5: Comparison results on GEdit-Bench. G_SC, G_PQ, and G_O refer to the metrics evaluated by GPT-4.1.
[…]
Instruct-P2P [3] 3.58 5.49 3.68
MagicBrush [56] 4.68 5.66 4.52
AnyEdit [49] 3.18 5.82 3.21
OmniGen [45] 5.96 5.89 5.06
Step1X-Edit [26] 7.09 6.76 6.70
BAGEL [9] 7.36 6.83 6.52
Gemini 2.0 [13] 6.73 6.61 6.32
GPT-4o [30] 7.85 7.62 7.53
UniWorld-V1 4.93 7.43 4.85
To validate out-of-domain generalization, we use GEdit-Bench [27] and report scores on the English dataset as shown in Table 5. UniWorld-V1’s perceptual score on GEdit-Bench (G_PQ) is relatively high (7.43), whereas its instruction-following score (G_SC) is relatively low (4.93). This is because most of our editing data originates from ImgEdit [48] and is limited in quantity, resulting in lower instruction diversity than BAGEL or Step1X-Edit. Additionally, text editing is a key evaluation metric in GEdit-Bench. Our dataset contains no text-editing samples, so it performs poorly in this aspect. Moreover, text editing is challenging for a 512-resolution SigLIP. We believe instruction-following capability can be improved by incorporating more data and fine-tuning the VLM. Regarding text-editing capability, we continue collecting relevant data and raising SigLIP’s resolution. Finally, we reiterate that our core contribution is finding that high-resolution SigLIP enables control over reference-image consistency.